the perimeter of a rectangular garden is 38 ft. the width of the gardem is one less than 3 times the length. determine the length and width of the garden

Answer by tommyt3rd(5050)

P=2(l+w)=38
1) l+w=19
2 w=3l-1

l+3l-1=19
4l=20
l=5
w=3*5-1=14